OpenStudy (linaminas):

If g (x) = 3x -8 find g(-1)

TheSmartOne (thesmartone):

plug in -1 in place of x :)

OpenStudy (rolldextro):

g(-1) just means evaluating the function at said value. as an example. f(x)=3x-8, evaluated at f(0) means f(0)=3(0)-8=-8. Now the same applies to your case.

OpenStudy (linaminas):

So g (-1) = 3 (-1) -8 ?

TheSmartOne (thesmartone):

Exactly!

OpenStudy (linaminas):

What's the next step to follow?

TheSmartOne (thesmartone):

What is 3* -1 = ? and then you subtract 8 from it and that's all :)

OpenStudy (linaminas):

Okay -11 thank you!

TheSmartOne (thesmartone):

Correct! Good job! :)

OpenStudy (linaminas):

Thank you (:

TheSmartOne (thesmartone):

It was my pleasure! :)